# Uncovering Bias Paths with LLM-guided Causal Discovery: An Active Learning and Dynamic Scoring Approach

### v1

@@ Abstract
Causal discovery (CD) plays a pivotal role in understanding the mechanisms underlying complex systems. While recent algorithms can detect spurious associations and latent confounding, many struggle to recover fairness-relevant pathways in realistic, noisy settings. Large Language Models (LLMs), with their access to broad semantic knowledge, offer a promising complement to statistical CD approaches, particularly in domains where metadata provides meaningful relational cues. Ensuring fairness in machine learning requires understanding how sensitive attributes causally influence outcomes, yet CD methods often introduce spurious or biased pathways. We propose a hybrid LLM-based framework for CD that extends a breadth-first search (BFS) strategy with active learning and dynamic scoring. Variable pairs are prioritized for LLM-based querying using a composite score based on mutual information, partial correlation, and LLM confidence, improving discovery efficiency and robustness.
To evaluate fairness sensitivity, we construct a semi-synthetic benchmark from the UCI Adult dataset, embedding a domain-informed causal graph with injected noise, label corruption, and latent confounding. We assess how well CD methods recover both global structure and fairness-critical paths (e.g., sex $\rightarrow$ education $\rightarrow$ income).
Our results show that LLM-guided methods, including the proposed method, demonstrate competitive or superior performance in recovering such pathways under noisy conditions. We highlight when dynamic scoring and active querying are most beneficial and discuss implications for fairness auditing in real-world datasets.
@@ 1 Introduction
Causal discovery (CD) identifies causal relationships among variables and is widely used in domains such as epidemiology, genetics, and economics (Kampani et al. 2024). Recent advances in CD algorithms have opened new avenues for addressing bias and fairness in machine learning (ML). Bias in ML often arises from hidden confounders or unequal treatment of sensitive groups (Barocas, Hardt, and Narayanan 2023). For instance, a hiring algorithm may inherit biased patterns from historical data that unfairly disadvantage certain demographic groups (Binns 2018). CD helps identify such pathways, distinguishing genuine effects from biases introduced by confounders and ensuring decisions are based on causal factors rather than proxies (Makhlouf, Zhioua, and Palamidessi 2020). Unlike traditional methods relying on statistical correlations or predefined fairness metrics, CD disentangles direct effects, indirect effects, and spurious correlations (Miguel-Hurtado et al. 2016).
This has prompted growing interest in using CD to support fairness-aware ML. Recent studies have proposed using counterfactual reasoning, structural causal models (SCMs), and fairness-specific constraints to detect and mitigate discriminatory pathways (Binkytė et al. 2023; Zuo et al. 2022; Wang and Luo 2023). However, classical CD methods often rely on strong assumptions, perform poorly under noise or confounding, and struggle to align with domain knowledge not explicitly encoded in data (Takayama et al. 2024).
Large Language Models (LLMs) have emerged as promising tools for CD, using their vast semantic knowledge to infer causal directions between variables such as “sex” and “income” from textual metadata (Le, Xia, and Chen 2024; Vashishtha et al. 2023; Khatibi et al. 2024; Kampani et al. 2024; Long, Schuster, and Piché 2023; Kıcıman et al. 2023). Jiralerspong et al. (Jiralerspong et al. 2024) proposed a breadth-first search (BFS) strategy for efficient graph construction via LLM queries, showing improved performance over random or exhaustive querying. However, a growing concern is that CD methods, especially when used in fairness audits, may introduce or amplify bias by inferring spurious links involving sensitive features. In fairness-sensitive contexts, incorrect causal attributions can lead to unjustified conclusions about discrimination or fairness violations. Thus, ensuring both the accuracy and interpretability of discovered pathways is essential for responsible ML deployment.
@@
Building on prior work, we introduce an enhanced BFS framework that integrates Active Learning (AL) and Dynamic Scoring. Instead of querying all pairs uniformly, our method ranks pairs using mutual information (MI), partial correlation (PCorr), and LLM-derived confidence, while discouraging redundant queries and focusing on informative regions via a history-based weight. This approach treats the LLM not as a passive responder but as an adaptive component within an iterative learning loop.
@@
A key challenge in evaluating CD methods in fairness contexts is the absence of real-world datasets with known ground-truth causal structures involving sensitive attributes (Loftus et al. 2018). Most, like the UCI Adult dataset, offer rich features but lack verified causal graphs. This limitation makes it difficult to validate whether inferred pathways truly reflect bias mechanisms or are artifacts of the discovery process. To address this, we construct a semi-synthetic benchmark that retains real-world semantics while embedding a known fairness-relevant graph for controlled evaluation.
We analyze the performance of our proposed method and other baseline CD methods on the constructed semi-synthetic benchmark based on the UCI Adult dataset that incorporates fairness-relevant structure, noise, and latent confounding. We evaluate these methods across their ability to recover key pathways (e.g., sex $\rightarrow$ education $\rightarrow$ income) and to support fairness audits.
@@
We also conduct a hyperparameter sensitivity analysis to examine how LLM-specific factors such as temperature, query prioritization weights, and stopping criteria affect CD performance. These findings offer practical insights into how LLM behavior and prompting influence graph reconstruction accuracy and fairness-path recovery.
Our contributions are as follows:
  • We propose a hybrid LLM-based CD method that integrates AL and dynamic scoring to prioritize informative queries during BFS-based graph recovery.
  • We introduce a semi-synthetic benchmark grounded in fairness literature, with latent confounding, label noise, and sensitive pathways, to evaluate CD methods under realistic conditions.
  • We provide a comparative analysis of pathway recovery across methods and examine how different discovery strategies affect fairness-related inferences.
  • We analyze the impact of LLM-specific hyperparameters on discovery quality.
@@ 2 Related Work
Causal discovery has been explored through statistical, optimization-based, and more recently, LLM-guided approaches (Long et al. 2025).
@@
Classical methods like the Peter-Clark (PC) algorithm (Spirtes and Glymour 1991) and Greedy Equivalence Search (GES) (Meek 1997) use conditional independence tests or score-based search but rely on strong assumptions (e.g., causal sufficiency, faithfulness) and degrade with latent confounding or limited data. Optimization-based approaches, such as NOTEARS (Zheng et al. 2018) and DAGMA (Bello, Aragam, and Ravikumar 2022), frame CD as a constrained optimization problem. While accurate, they are computationally expensive and sensitive to hyperparameters, limiting their use in fairness-critical or iterative settings.
LLM-based CD methods offer an alternative by leveraging variable descriptions to infer structure, offering a way to encode human knowledge and priors. Kıcıman et al. (Kıcıman et al. 2023) introduced pairwise LLM queries using only metadata, though the quadratic query complexity limits scalability. Others combine LLM-derived causal orders with algorithms like PC (Vashishtha et al. 2023), or embed LLMs in multi-agent or prompt-refinement loops (Khatibi et al. 2024; Le, Xia, and Chen 2024). Kampani et al. (Kampani et al. 2024) take a hybrid approach, using LLMs to generate priors, refining them with methods like NOTEARS.
Takayama et al. (Takayama et al. 2024) propose statistical causal prompting, where LLMs enhance statistical methods by providing priors or plausible causal orders. While robust, this one-time augmentation lacks iterative refinement. Similarly, Jiralerspong et al. (Jiralerspong et al. 2024) propose a BFS framework to reduce LLM query complexity during structure discovery. Their approach treats LLM responses as inputs for efficient edge discovery, dramatically lowering runtime without sacrificing metadata grounding. However, BFS applies fixed selection heuristics and does not dynamically balance statistical evidence with LLM insights, making it susceptible to early decision errors.
Our method builds on this BFS foundation, introducing AL and a dynamic scoring strategy that prioritizes variable pairs based on a mixture of MI, PCorr, and LLM confidence scores. Compared to prior methods that treat all variable pairs uniformly or rely solely on LLMs or statistics, our approach dynamically balances complementary signals and avoids over-reliance on any single information source.
Some of the above-mentioned approaches differ fundamentally from our framework in both scope and evaluation goals. Methods such as those proposed by Vashishtha et al. (Vashishtha et al. 2023), Khatibi et al. (Khatibi et al. 2024), Le et al. (Le, Xia, and Chen 2024), and Kampani et al. (Kampani et al. 2024) use LLMs to generate global causal orders or priors that are later refined using constraint-based or optimization-based algorithms like PC or NOTEARS. In contrast, our method treats the LLM as an active querying agent embedded in an iterative loop, where causal judgments are elicited at the pairwise level and guided dynamically by a scoring function informed by both statistical and historical signals. Unlike prompting-based or prior-initialized approaches, our framework does not rely on a static, up-front extraction of priors, but instead continuously adapts query selection through AL.
We further emphasize fairness-aware causal recovery, targeting sensitive paths and mediators (e.g., sex $\rightarrow$ education $\rightarrow$ income). This responds to calls for causal approaches to fairness (Nabi and Shpitser 2018; Kilbertus et al. 2017; Zhang and Bareinboim 2018; Chiappa 2019), highlighting the need for explainable, intervenable, and robust fairness assessments under noisy conditions. In this sense, our method contributes to a growing understanding of how LLM-based querying, when combined with statistical reasoning, can affect fairness-relevant causal inference under noisy and complex conditions.
@@ 3 Proposed LLM-Based Method
3.1 LLM-Guided Breadth-First Search for Causal Discovery
The foundation of the proposed method is the BFS-based CD framework introduced by (Jiralerspong et al. 2024). This approach iteratively queries an LLM to infer causal relationships among variables in a data set and reduces the number of queries required since visiting each node once is sufficient to determine all edges originating from that node. Using this prompting approach, each query is approached as an opportunity for node expansion within a BFS algorithm, progressively building the causal graph as it is traversed using BFS. To effectively implement the BFS, it is essential to determine the sequence in which nodes are explored. For this purpose, the Directed Acyclic Graph (DAG) nature of causal graphs is leveraged, by using their topological ordering to guide the BFS traversal. The BFS process follows three main steps:
  1. 1.
    Initialization: The algorithm begins by querying the LLM to identify independent variables that are not caused by any other variables which form the starting points for graph exploration.
  2. 2.
    Expansion: For each variable on the current frontier (independent variables or those already explored), the LLM is queried about which other variables it causally influences, using prompts that incorporate discovered structure so far.
  3. 3.
    Graph Construction: Inferred edges are added to a directed graph if they do not introduce cycles. Each new relationship is checked to ensure that the variables in the frontier are connected to previously unexplored nodes. The process continues until all nodes are explored. The final output is a predicted causal graph in the form of an adjacency matrix.
While more efficient than exhaustive pairwise querying, this approach still explores variable pairs uniformly. As a result, many LLM queries target uninformative pairs, especially in sparse or noisy graphs, motivating the need for adaptive query selection.
@@ 3.2 Active Learning and Dynamic Scoring for Efficient Querying
To improve the informativeness of queries and reduce redundancy, we introduce a scoring-based AL framework. This component augments the BFS approach with a dynamic prioritization mechanism for selecting the most valuable variable pairs to query. Figure 1 illustrates the full framework.
@@
For each unqueried pair $(x,y)$, a dynamic informativeness score $S(x,y)$ is computed as:
@@
  • Statistical Score: Combines mutual information (MI) and partial correlation (PCorr), measuring both linear and non-linear dependencies:
@@
Query Selection via Active Learning
We incorporate dynamic scoring into an AL loop that determines which pairs to query. The query process iteratively selects:
$(x^{*},y^{*})=\arg\max_{(x,y)\in\text{UnqueriedPairs}}S(x,y)$ (5)
An LLM is then prompted to assess whether $x^{*}$ causally affects $y^{*}$ using the natural language format described in listing 1.
Listing 1 Causal query prompt format
Does X cause Y? Provide the answer in the format: <Answer>Yes</Answer> or <Answer>No</Answer>.
If the LLM predicts a positive causal link and the resulting graph remains acyclic, the edge $x^{*}\rightarrow y^{*}$ is added; otherwise, it is discarded. To estimate the LLM uncertainty, we extract token-level log-probabilities and compute the average probability of the model’s response (“Yes” or “No”). This score, scaled to [0,1], informs the dynamic scoring function. If log-probs are unavailable, a default confidence of 0.5 is used. The process terminates after a query limit or when scores fall below a threshold.
Cycle Detection and Output Construction
@@
4 Fairness Evaluation through Causal Pathway Analysis
To evaluate the fairness utility of learned causal graphs, we perform pathway analysis to quantify how sensitive variables influence target outcomes. Unlike classical fairness metrics based on statistical associations (e.g., demographic parity), pathway analysis distinguishes between direct, mediated, and spurious effects, offering a principled causal perspective to fairness assessment (Pearl 2009, 2022).
We analyze the structure and influence of paths from sensitive attributes $S$ (e.g., race, sex) to outcome variables $Y$ (e.g., income) in the learned graphs. This approach distinguishes between direct, mediated, and spurious influences, going beyond traditional statistical fairness metrics that lack causal grounding (Pearl 2022).
Path Count Comparison.
For each discovered graph, we enumerate and classify all directed paths originating from sensitive variables:
  • Direct Paths ($P_{\text{direct, fairness}}$): Number of direct causal edges ($S\rightarrow Y$) that connect the defined sensitive variables to the outcomes.
  • Indirect Paths ($P_{\text{indirect, fairness}}$): Number of causal paths ($S\xrightarrow{}\text{intermediate nodes}\xrightarrow{}Y$) in which sensitive variables influence the outcomes indirectly via intermediate variables.
  • Spurious Paths ($P_{\text{spurious, fairness}}$): Number of paths that involve sensitive variables but do not influence the outcomes ($S\xrightarrow{}\text{non-$Y$ nodes}$).
To assess structural alignment with fairness-relevant pathways, we report the number of direct and indirect paths for each graph and compare these across CD methods and to the known ground-truth graph. A higher overlap with true fairness-relevant paths indicates better recovery of bias mechanisms.
Quantitative Effect Decomposition.
To assess the magnitude of fairness-related effects, we estimate:
  • Direct Effect (DE): The estimated causal effect of $S$ on $Y$ along direct paths.
  • Indirect Effect (IE): The effect of $S$ on $Y$ through mediating variables.
  • Total Effect (TE): The sum of direct and indirect effects:
@@ 4 Fairness Evaluation through Causal Pathway Analysis
These effects are computed using structural equations or interventional estimators (depending on the graph), and normalized to account for scale:
@@
where $C_{\text{bias}}$ represents the total fairness-relevant contribution of $S$ to variance in the outcome. This enables comparative evaluation across datasets and methods. It quantifies the total effect of sensitive attributes on the outcome (income), scaled by outcome variance. A higher $C_{\text{bias}}$ indicates a greater risk of propagating sensitive information through direct or mediated pathways, particularly important for fairness auditing.
5 Synthetic Graph Construction Based on UCI Adult Attributes
To evaluate fairness-aware CD under controlled yet realistic conditions, we construct a semi-synthetic dataset based on the UCI Adult dataset (Kohavi et al. 1996), commonly used in fairness research for its demographic diversity and income disparities. Our synthetic graph retains the original 15 variables and encodes plausible causal links informed by prior work (Kilbertus et al. 2017; Nabi and Shpitser 2018).
The DAG assumes that demographic features (e.g., race, sex) influence education and occupation; education affects income both directly and via mediators like occupation and capital gains; and age and marital status serve as confounders or mediators in employment-related paths. Capital-related features are modeled as downstream effects of occupational history, contributing to income.
To simulate structural unfairness, we inject direct edges from sex and race to income, consistent with causal fairness benchmarks (Nabi and Shpitser 2018; Kilbertus et al. 2017). These pathways are complemented by indirect paths (e.g., sex $\rightarrow$ education $\rightarrow$ occupation $\rightarrow$ income) to facilitate the estimation of direct, indirect, and total effects in fairness evaluation. The baseline graph and technical details on the data generation process can be found in Appendix A.1.
5.1 Structured Noise Augmentation
Additive Noise for Continuous Variables.
For each continuous variable X, noise is sampled as:
Symmetric Label Noise in Categorical Variables
$P(\tilde{Y}_{i}=y^{\prime})=\begin{cases}1-\gamma&\text{if }y^{\prime}=Y_{i}\\ \frac{\gamma}{K-1}&\text{if }y^{\prime}\neq Y_{i}\end{cases}$ (9)
This means with probability $\gamma$, the label is flipped uniformly to one of the $K-1$ incorrect categories. And with probability $1-\gamma$, the label remains correct. This uniform corruption model known as symmetric label noise, is widely used in label-noise robust learning (Natarajan et al. 2013; Patrini et al. 2017).
Latent Confounding via Unobserved Variable $U$
To simulate unmeasured social influences (e.g., privilege or institutional access), we introduce a latent confounder $U\sim\mathcal{N}(0,1)$ that affects intermediate variables such as years of education, capital gain, and work hours, thereby indirectly influencing income. This reflects latent societal advantages that are not captured in observed features.
The latent confounder is injected via additive linear terms into structural equations following standard assumptions in structural causal models (SCMs). For the binary outcome income, we use a nonlinear threshold function that combines observable and latent contributions. This setup reflects bias amplification near decision boundaries and is consistent with prior fairness simulation frameworks. Full equations and implementation details are provided in Appendix A.2.
Controlled Replication Across Seeds
To evaluate generalization and fairness robustness, we generate four datasets using different random seeds, each with the same underlying graph but distinct noise realizations. Noise is added post-graph generation, preserving causal structure while introducing realistic variability. This mirrors standard practices in fairness benchmarking and robustness testing for CD.

### v2

@@ Abstract
Ensuring fairness in machine learning requires understanding how sensitive attributes like race or gender causally influence outcomes. Existing causal discovery (CD) methods often struggle to recover fairness-relevant pathways in the presence of noise, confounding, or data corruption. Large language models (LLMs) offer a complementary signal by leveraging semantic priors from variable metadata. We propose a hybrid LLM-guided CD framework that extends a breadth-first search strategy with active learning and dynamic scoring. Variable pairs are prioritized for querying using a composite score combining mutual information, partial correlation, and LLM confidence, enabling more efficient and robust structure discovery. To evaluate fairness sensitivity, we introduce a semi-synthetic benchmark based on the UCI Adult dataset, embedding domain-informed bias pathways alongside noise and latent confounders. We assess how well CD methods recover both global graph structure and fairness-critical paths (e.g., sex→education→income). Our results demonstrate that LLM-guided methods, including our active, dynamically scored variant, outperform baselines in recovering fairness-relevant structure under noisy conditions. We analyze when LLM-driven insights complement statistical dependencies and discuss implications for fairness auditing in high-stakes domains.
@@ 1 Introduction
Bias in machine learning (ML) systems affects decisions in hiring, lending, education, and healthcare. These biases often emerge through indirect, structural pathways where sensitive attributes (e.g., race or gender) influence outcomes via proxies or confounded relationships (Graetz, Boen, and Esposito 2022). Standard fairness audits rely on statistical disparity metrics, but such metrics overlook how bias propagates causally (Chinta et al. 2025). Without structural insight, interventions may be ineffective or misleading.
Causal discovery (CD) offers tools to identify such pathways, distinguishing genuine effects from those introduced by confounders or measurement artifacts. While recent methods apply structural causal models or fairness constraints to audit ML systems, classical CD techniques often fail under noisy data, latent confounding, or incomplete metadata. These limitations are acute in fairness contexts where both precision and interpretability are critical (Takayama et al. 2024).
Large language models (LLMs) have emerged as promising tools for CD, using their vast semantic knowledge to infer causal directions between variables from textual metadata (Le, Xia, and Chen 2024; Vashishtha et al. 2023). Prior work has used LLMs to infer causal links or orderings, and to reduce query cost via strategies like breadth-first search (BFS) (Jiralerspong et al. 2024). However, naive LLM use risks over or under attributing causality, especially involving sensitive attributes, raising concerns about spurious fairness conclusions.
@@
Building on prior work, we introduce a fairness-driven CD framework that enhances the BFS approach with active learning (AL) and dynamic scoring. Our method prioritizes variable pairs using mutual information (MI), partial correlation (PCorr), and LLM-derived confidence, while discounting redundant queries and focusing on informative regions via a history-based weight. Unlike exhaustive querying or fixed-order search, our adaptive strategy balances semantic and statistical cues to recover fairness-relevant paths more efficiently.
@@
A key challenge in evaluating CD methods in fairness contexts is the absence of real-world datasets with known ground-truth causal structures involving sensitive attributes (Loftus et al. 2018). Most, like the UCI Adult dataset, offer rich features but lack verified causal graphs. To enable robust evaluation, we construct a semi-synthetic benchmark from the UCI Adult dataset, embedding a known fairness-relevant causal graph with noise, corruption, and confounding. We assess how well CD methods recover both global structure and key fairness pathways (e.g., sex $\rightarrow$ education $\rightarrow$ income).
@@
We also conduct a hyperparameter sensitivity analysis to examine how LLM-specific factors such as temperature, query prioritization weights, and stopping criteria affect CD performance. These findings offer practical insights into how LLM behavior and prompting influence graph reconstruction accuracy and fairness-path recovery. Our approach is designed for robustness, interpretability, and social accountability, enabling stakeholders to trace how outcomes are structurally linked to sensitive variables and to intervene in meaningful ways.
Our key contributions:
  • A hybrid LLM-CD framework with active learning and dynamic scoring to enhance fairness-aware structure discovery.
  • A benchmark for fairness-sensitive CD grounded in real-world semantics and controlled ground truth.
  • An evaluation of pathway recovery across methods, highlighting conditions where LLM-guided discovery improves fairness diagnostics.
  • A sensitivity analysis on LLM parameters and their influence on recovery quality.
@@ 2 Related Work
Statistical and Optimization-Based CD. CD methods span statistical, optimization-based, and more recently, LLM-guided approaches (Long et al. 2025). Classical techniques such as PC (Spirtes and Glymour 1991) and GES (Meek 1997) rely on conditional independence tests and strong assumptions (e.g., faithfulness), often failing under latent confounding or limited data. Optimization-based methods like NOTEARS (Zheng et al. 2018) and DAGMA (Bello, Aragam, and Ravikumar 2022) offer greater precision but are computationally expensive and sensitive to hyperparameters, making them less practical for fairness-critical contexts.
LLM-Guided CD. Recent LLM-based approaches use variable semantics to infer structure, with Kıcıman et al. (Kıcıman et al. 2023) proposing pairwise causal queries from metadata. However, such methods scale poorly due to quadratic query complexity. Others combine LLM-inferred causal orders with classical CD (Vashishtha et al. 2023), or use multi-agent prompting strategies (Khatibi et al. 2024; Le, Xia, and Chen 2024). Kampani et al. (Kampani et al. 2024) generate LLM-based priors refined through NOTEARS. Takayama et al. (Takayama et al. 2024) introduced statistical causal prompting, where LLMs enhance statistical methods by providing priors or plausible causal orders. Jiralerspong et al. (Jiralerspong et al. 2024) proposed a BFS strategy to reduce query complexity, but rely on fixed pair ordering, making them susceptible to early decision errors.
We build on this line by introducing AL and a dynamic scoring mechanism to adaptively prioritize informative variable pairs using MI, PCorr, and LLM confidence. Our method dynamically balances semantic and statistical signals, in contrast to prior work that uses LLMs for one-time priors or uniform querying. We focus on observational CD methods to maintain comparability with LLM-based approaches, leaving integration with interventional or experimental methods (e.g., do-calculus, IV-based CD) for future work.
Fairness and Causal Pathways. Beyond structural accuracy, recent work has emphasized the importance of CD for fairness auditing. Causal fairness frameworks (Nabi and Shpitser 2018; Kilbertus et al. 2017; Zhang and Bareinboim 2018; Chiappa 2019) focus on identifying discriminatory pathways, e.g., sex $\rightarrow$ education $\rightarrow$ income that explain how bias propagates through mediation or proxy variables. Our method is explicitly tuned to recover such fairness-relevant paths with greater robustness under noise and limited samples, enabling more actionable fairness assessments.
Social and Regulatory Context. Beyond computer science, our work connects to social science perspectives. Structural discrimination and intersectionality highlight how compounding disadvantage flows through proxy variables, not just direct links (Crenshaw 2013; Bonilla-Silva 1997). Regulatory frameworks such as GDPR’s “right to explanation” and EEOC compliance requirements also motivate interpretable, path-level audits (Kaminski 2021).
Causal Inference and Accountability. We also draw from causal inference frameworks like the potential outcomes model (Rubin 1974; Imbens and Rubin 2015) and matching-based fairness analysis (Stuart 2010), which stress structural validity and transparency of assumptions. Our hybrid design reflects these values by surfacing interpretable causal hypotheses grounded in both observed data and language-based priors. Finally, we respond to calls for auditable AI systems through transparency-centered work on model documentation and sociotechnical accountability (Mitchell et al. 2019; Selbst et al. 2019).
@@ 3 Proposed LLM-Based Method
We build on the BFS-based CD framework of (Jiralerspong et al. 2024), where an LLM is queried iteratively to construct a causal graph. The method begins by identifying independent root variables and expands outward via LLM-guided queries, treating each confirmed edge as a BFS traversal step. To ensure acyclicity, edges are added only if they do not create cycles. While more efficient than exhaustive querying, this approach treats all variable pairs uniformly and can waste queries on uninformative or redundant relationships.
@@ 3.1 Active Learning and Dynamic Scoring for Efficient Querying
We introduce an AL strategy that selects variable pairs based on a dynamic informativeness score. This score integrates statistical signals, model confidence, and query history, prioritizing pairs likely to yield informative causal judgments. The query loop continues until a maximum iteration or informativeness threshold is reached. Figure 1 summarizes the overall architecture.
@@
Each unqueried pair $(x,y)$ receives a composite score:
@@
  • Statistical Score: Combines MI and PCorr, measuring both linear and non-linear dependencies:
@@
    PCorr is computed conditionally on the current discovered parent sets, following the iterative conditioning procedure of the PC algorithm.
@@
Querying the LLM
At each step, the pair $(x^{*},y^{*})$ with the highest score is selected:
$(x^{*},y^{*})=\arg\max_{(x,y)\in\text{Unqueried}}S(x,y)$ (5)
The LLM is queried using a simple natural language prompt:
Listing 1 Simplified LLM Query Format
1System: You are a domain expert.
2User: You assist the expert in evaluating possible causal links.
3
4Metadata: Each variable is described using a short name and a natural language description.
5
6Query: Does VAR_A cause VAR_B?
7Respond with <Answer>Yes</Answer> or <Answer>No</Answer>.
Domain-specific instructions and variable descriptions are included to contextualize each query. See Appendix for full example prompts.
If the LLM returns “Yes” and adding $x^{*}\rightarrow y^{*}$ does not form a cycle, the edge is added to the graph. To estimate the LLM uncertainty, we extract token-level log-probabilities and compute the average probability of the model’s response (“Yes” or “No”). This score, scaled to [0,1], informs the dynamic scoring function. If log-probs are unavailable, a default confidence of 0.5 is used. The process terminates after a query limit or when scores fall below a threshold.
@@
4 Fairness Evaluation via Pathway Analysis
We assess the fairness utility of learned causal graphs by analyzing how sensitive attributes $S$ (e.g., race, sex) influence outcomes $Y$ (e.g., income). Unlike statistical fairness metrics, pathway analysis provides a causal perspective by distinguishing direct, mediated, and spurious effects (Pearl 2009, 2022).
Path Classification
We enumerate all directed paths from $S$ and classify them as:
  • Direct: Edges $S\rightarrow Y$.
  • Indirect: Paths $S\rightarrow\cdots\rightarrow Y$ via mediators.
  • Spurious: Paths involving $S$ that do not reach $Y$.
Comparing these across methods and against the ground-truth graph reveals how well fairness-relevant mechanisms are recovered.
Effect Decomposition
We estimate the causal effects of $S$ on $Y$ using either structural equations or interventional estimators:
$DE$ $:\text{Direct effect (via }S\rightarrow Y\text{)}$
$IE$ $:\text{Indirect effect (via mediators)}$
@@ 4 Fairness Evaluation via Pathway Analysis
We normalize by outcome variance to obtain:
@@
Here, $C_{\text{bias}}$ quantifies the fairness-relevant contribution of $S$ to $Y$, enabling method comparison across datasets. Higher values indicate a greater risk of bias propagation.